When 112.4 g of barium nitrate react with 220 g of aluminum chloride, how many grams of each product can be produced? What is the limiting reactant?

 


1
Expert's answer
2021-02-12T05:03:53-0500

Mass of barium nitrate= 112.4g

Molar mass of barium nitrate= 261

Mass of aluminium chloride=220g

Molar mass of aluminium chloride= 133g/mol

Balance equation is =

3Ba(NO3)2 +2AlCl3= 3BaCl2+ 2Al(NO3)3

Mole of barium nitrate= mass/molar mass

Mole= 112.4÷261=0.43

Mole of aluminium chloride= 1.65

So here is mole of barium nitrate is less so it will be limiting reagent.